### v2

# Enhanced Robot Arm at the Edge with NLP and Vision Systems

Abstract
This paper introduces a ”proof of concept” for a new approach to assistive robotics, integrating edge computing with Natural Language Processing (NLP) and computer vision to enhance the interaction between humans and robotic systems. Our ”proof of concept” demonstrates the feasibility of using large language models (LLMs) and vision systems in tandem for interpreting and executing complex commands conveyed through natural language. This integration aims to improve the intuitiveness and accessibility of assistive robotic systems, making them more adaptable to the nuanced needs of users with disabilities. By leveraging the capabilities of edge computing, our system has the potential to minimize latency and support offline capability, enhancing the autonomy and responsiveness of assistive robots. Experimental results from our implementation on a robotic arm show promising outcomes in terms of accurate intent interpretation and object manipulation based on verbal commands. This research lays the groundwork for future developments in assistive robotics, focusing on creating highly responsive, user-centric systems that can significantly improve the quality of life for individuals with disabilities.
I INTRODUCTION
Figure 1: The architecture of the proposed system. Initially, a user’s spoken input is captured and processed by VOSK Speech to Text. This text is then fed into GPT API, which interprets the text and generates a response in JSON format. Subsequently, this JSON output is translated into a ROS (Robot Operating System) Command, which acts as a bridge between high-level commands and robotic control. Finally, the ROS Command triggers the desired Robot Action, completing the process from human speech to robotic operation.
The advent of assistive robotics [1] has marked a significant milestone in leveraging technology for augmenting human capabilities, especially for individuals with disabilities [2]. These robotic systems offer an unprecedented potential to assist in daily activities [3, 4], thereby enhancing the quality of life and fostering greater independence for people facing physical challenges. The integration of robotics in assistive technologies ranges from sophisticated prosthetic limbs [5] to autonomous wheelchairs and robotic arms [6] designed for home and workplace environments. The primary objective behind these innovations is to bridge the gap between human intentions and the physical execution of tasks, making everyday activities more accessible to those with limited mobility or dexterity.
Despite considerable advancements in mechanical design and sensor technology, one of the persistent challenges in assistive robotics is developing intuitive and natural interaction interfaces [7] based on the specific needs of the users. Traditional control mechanisms, such as joysticks [8], or touch screens [9], often require a degree of fine motor control that might not be feasible for all users. Furthermore, these control schemes can be cumbersome and unintuitive, particularly for complex tasks that involve multiple steps or precise movements [10]. As an example, users of the assistive robotic arms mounted on a powered wheelchair must alternate between various control settings to operate a 7-degree-of-freedom (DOF) arm and its preset positions using a 2-axis wheelchair joystick [11, 12]. Voice control is also implemented in some assistive robotic arms mounted on wheelchairs but user studies conducted by Chung et. al. [13] showed that voice control, while beneficial for hands-free operation, often struggles with accuracy in noisy environments and can be slow for executing a series of commands, leading to frustration among users. These limitations underscore the need for more adaptive and user-friendly methods of communication between humans and assistive robotic systems.
Natural Language Processing (NLP) emerges as a promising solution to this challenge, offering a more accessible and intuitive means for users to interact with robotic systems [14]. The rise of large language models (LLMs) like GPT (Generative Pre-trained Transformer) [15] further enhances the potential of NLP in robotics [16, 17]. These models, trained on vast amounts of text data, can understand and generate human-like text, allowing for more nuanced and complex interactions between humans and robots. Different research groups have started to use LLMs to control robots and autonomous vehicles [18]. Despite these advancements, research [19] showed that one of the main challenges in assistive robotics is the translation of natural language into robot actions that are both precise and contextually appropriate. On the other hand, the specific needs of users have often not been the central focus in the design of NLP-powered robots. There remains a significant opportunity to tailor these technologies to better serve individuals with diverse requirements, particularly in assistive applications. By prioritizing user-centric development, future research can bridge the gap between technological capability and practical usability, ensuring that assistive robots are not only advanced in their understanding and execution of natural language commands but also finely attuned to the nuanced needs and preferences of their human users.
Edge computing [20, 21] represents a significant advancement in data processing for robotics [22, 23], particularly in the realm of assistive technologies. By localizing computation to data sources—such as robots and sensors—it not only minimizes latency, essential for the safety and efficacy of assistive robots responding to instant commands, but also enables these systems to operate independently of internet connections, enhancing autonomy and safety that research [24] has shown are important factors in designing assistive technology. The integration of edge computing with advanced technologies like LLM and NLP facilitates the creation of highly responsive, personalized robotic systems capable of understanding and executing complex natural language commands in real-time. This approach not only meets the specific needs and preferences of users, making assistive devices more intuitive and effective for individuals with disabilities but also propels the field towards more adaptive and autonomous solutions.
This paper serves as a ”proof of concept” that combines edge computing with LLM, NLP, and computer vision to control robots with the hope of eventually advancing the field of assistive robotics towards highly responsive and user-centric assistive robotic systems. In this paper, we have implemented our proposed method on a robotic arm as a ”proof of concept” and current ongoing research involves implementing our method on a wheelchair-mounted assistive robot arm (Gen3 robot arm) and collaborating with Kinova, a leading assistive robotics company, and a neuromuscular neurologist to conduct user studies to tailor and enhance our technology based on direct feedback, targeting the unique needs and preferences of users. Emphasizing a user-centric design philosophy is crucial in our project to develop assistive robots that significantly improve the quality of life for individuals with disabilities, leading to more personalized and impactful robotic aids.
II Methods
II-A Architecture
Figure 1 shows the architecture of the proposed system where spoken language is converted to robot actions. Speech is first transcribed into text, then interpreted by GPT API into JSON formatted commands. These commands are translated into ROS (Robot Operating System) instructions, which ultimately drive the robot to perform the corresponding actions.
II-B Experimental Setup
We used 4 Degrees of Freedom (DOF) PincherX 100 robotic arm and the vision kit (Intel D415 RealSense Depth Camera, which features $1920\times 1080$ resolution at 30fps with a $65^{o}\times 40^{o}$ field of view and an ideal range of 0.5m to 3m) from Trossen Robotics, IL, USA for our poof-of-concept experiments. This robot arm is controlled by Robot Operating System (ROS2) Humble running on Ubuntu 22.04 operating system. The computer used is a Dell Precision 3660 Tower Core i7 with 32GB DDR5 up to 4400MHz UDIMM non-ECC memory and Nvidia RTXA4000 GPU. For audio capture, a Blue Yeti X USB Microphone is used to ensure high-quality input for the speech-to-text conversion process. Figure 2 shows the experimental setup.
Figure 2: For our experiments, we used a 4 DOF PincherX 100 robotic arm and an Intel D415 RealSense Depth Camera from Trossen Robotics, IL, USA. The robot arm, controlled via ROS2 Humble on Ubuntu 22.04, was connected to a Dell Precision 3660 with an i7 processor, 32GB RAM, and Nvidia RTXA4000 GPU. Audio is captured using a Blue Yeti X USB Microphone.
II-C Screw Theory-based Numerical Inverse Kinematics of the Robot Arm Using Newton-Raphson Iterative Method
Inverse kinematics is used to map the desired position of the end-effector of the robot arm to joint positions. We have used numerical inverse kinematics based on Screw Theory in Robotics [25, 26] using Newton-Raphson iterative method. The following algorithm is implemented to find the numerical inverse kinematics of our robot arm:
  1. 1.
    Initialization: Given $T_{sd}$ (transformation of the desired frame with respect to space frame) and an initial guess of joint variables $q_{o}\in\mathbb{R}^{n}$, set $i=0$.
  2. 2.
    Set $[{\mathcal{V}}_{b}]=log(T_{sb}^{-1}(q_{i})T_{sd})$. While the algorithm is not converged:
    • Set $q_{i+1}=q_{i}+J^{\dagger}_{b}(q_{i}){\mathcal{V}}_{b}$.
    • Increment $i$,
where $q_{i}$ is the current joint angles guess, $\mathcal{V}_{b}$ is the body twist calculated from the error between the current end-effector configuration and the desired configuration, and $J_{b}^{\dagger}(q_{i})$ is the pseudoinverse of the body Jacobian matrix evaluated at the current joint angles $q_{i}$ [27].
For our robot arm that we use for the experiments, the screw axis of each joint expressed in the base frame in the robot’s zero position (see Figure 3 for link lengths and screw axes assignments) are as follows:
$\mathcal{S}_{1}$ $={\begin{bmatrix}0&0&1&0&0&0\end{bmatrix}}^{T},$ (1)
$\mathcal{S}_{2}$ $={\begin{bmatrix}0&1&0&-0.08945&0&0\end{bmatrix}}^{T},$
$\mathcal{S}_{3}$ $={\begin{bmatrix}0&1&0&-0.18945&0&0.035\end{bmatrix}}^{T},$
$\mathcal{S}_{4}$ $={\begin{bmatrix}0&1&0&-0.18945&0&0.135\end{bmatrix}}^{T},$
Figure 3: Robot arm depicted in home position with screw axes and link lengths assignments that are essential to finding the forward kinematics. The lengths depicted on the bottom photo are: $L_{1}=0.08945$, $L_{2}=0.1$, $L_{m}=0.035$, $L_{3}=0.1$, and $L_{4}=0.08605$
and thus, $T_{sb}(q)$ (the pose of the end-effector frame $\{b\}$ in the base frame $\{s\}$) can be written as the following equation using the Product of Exponentials PoE formula:
$T_{sb}(q)=e^{[\mathcal{S}_{1}]q_{1}}e^{[\mathcal{S}_{2}]q_{2}}e^{[\mathcal{S}_{3}]q_{3}}e^{[\mathcal{S}_{4}]q_{4}}M,$ (2)
where, $q_{1},q_{2},q_{3},\text{ and }q_{4}$ are joint angles, and $M\in SE(3)$ is the end-effector configuration when the robot is at its zero position and for our robot arm it is represented by the following matrix:
$M=\begin{pmatrix}1&0&0&0.22105\\ 0&1&0&0\\ 0&0&1&0.18945\\ 0&0&0&1\end{pmatrix}.$ (3)
The Space Jacobian of this robot, $J_{s}(q)$ can be expressed as:
$J_{s}(q)=\begin{bmatrix}J_{s1}&J_{s2}&J_{s3}&J_{s4}\end{bmatrix},$ (4)
where, $J_{s1}$, $J_{s2}$, $J_{s3}$, and $J_{s4}$ can be calculated using the following equations:
$\begin{array}[]{l}J_{s1}=\mathcal{S}_{1}\\ J_{s2}=[Ad_{e^{[\mathcal{S}_{1}]q_{1}}}]\mathcal{S}_{2}\\ J_{s3}=[Ad_{e^{[\mathcal{S}_{1}]q_{1}}e^{[\mathcal{S}_{2}]q_{2}}}]\mathcal{S}_{3}\\ J_{s4}=[Ad_{e^{[\mathcal{S}_{1}]q_{1}}e^{[\mathcal{S}_{2}]q_{2}}e^{[\mathcal{S}_{3}]q_{3}}}]\mathcal{S}_{4}\end{array}$ (5)
in which, $\mathcal{S}_{i}$ is an expression for the screw axis describing the ith joint axis in terms of the fixed frame with the robot in its zero position given above, and $[Ad_{T_{i-1}}]\mathcal{S}_{i}$ is the screw axis describing the ith joint axis, but after it undergoes the rigid body displacement $T_{i}$ instead of being at zero position. In other words, it is the Adjoint map of the screw axis for when the robot is no longer in zero position. Note that if $T=(R,p)\in SE(3)$ is a transformation matrix where $R$ is the rotation matrix and $p$ is the position vector, then its adjoint representation $[Ad_{T}]$ can be calculated by:
$[Ad_{T}]=\begin{pmatrix}R&o\\ [p]R&R\end{pmatrix}\in\mathbb{R}^{6\times 6}.$ (6)
Note that the bracket notation $[p]$ is the $3\times 3$ skew-symmetric matrix representation of the position vector $p$. The body Jacobian can then be calculated from the space Jacobian using the Adjoint transformation:
$J_{b}(q)=[Ad_{T_{bs}}]J_{s}(q),$ (7)
where, $T_{bs}=T^{-1}_{sb}$ is the inverse of the homogeneous matrix $T_{sb}$ that was derived from forward kinematics earlier.
II-D Vision‐Language-Aided Inverse Kinematics Control of the Robot Arm
II-D1 Vision-Aided Control
The objective here is to merge the numerical inverse kinematics with the robot’s perception package to create a vision-aided inverse kinematics motion planner. The robot’s perception pipeline uses a combination of depth data and color information from the RealSense camera to create a point cloud for 3D object detection and manipulation. The depth camera first captures the AprilTag [28] attached to the robot’s arm to help ROS figure out the homogeneous transformation of the robot’s base frame relative to the camera and vice versa. This helps convert the camera’s depth readings of scene objects to homogeneous transformations with respect to the robot’s base. Later, the inverse kinematics module maps these transformations to joint angle set-points to make the robot manipulate the objects. Stereo vision technology alongside the infrared sensor helped estimate the 3D structure of the scene.
II-D2 Speech-to-Text Integration
Our program incorporates an Open Source speech recognition program called VOSK by Alpha Cephei Inc. VOSK operates offline and runs entirely on the local machine. This is crucial for the program as it no longer requires voice activation to require internet connectivity to transcribe the speech. Once the speech is captured and processed by VOSK, the program takes over to analyze the transcribed text. This analysis involves parsing the text output to understand the commands and translating them into specific actions that control the robot.
II-D3 ChatGPT API Integration
OpenAI provides API access to their LLM and for this program, we used GPT-3.5 and later GPT-4 (for comparison with GPT-3.5 tests) as our models. The API was fine-tuned to comprehend and transcribe natural spoken language, significantly enhancing accuracy and context relevance for the process. The user’s speech that has previously been parsed by VOSK is now processed by the LLM. We leveraged the OpenAI API to develop a program that takes the text input and converts it into a JSON-formatted list of instructions. This output gave us certain key objectives such as the action, the object, the direction, and so forth. For example, when given a phrase such as ”it’s opposite day, now move to the right”, the API integration can understand key terms ”opposite” and ”right” leading the robot to move to the left; thus interpreting intent rather than solely reading ”right” and moving to the right as is common in voice control devices. This implementation produces an increase in language the robot can interpret, allowing the user to speak in a natural manner.
II-D4 Language-Vision-aided Control
In our approach to create a language-vision-aided control system for robotic manipulation, we integrated the LLM with advanced computer vision technologies. The method involved first processing verbal commands through the LLM to interpret the user’s intent accurately. These commands were then translated into specific parameters for the vision system, guiding the robot to identify and interact with objects based on their characteristics, such as color. Our experimental setup included objects of three distinct colors (red, blue, and green) placed within the robot’s workspace. The system’s objective was to accurately decode spoken commands like ”Pick up the red cube,” and through a seamless integration of NLP and computer vision, enable the robot to visually locate, identify, and manipulate the specified object. This method aimed to test the robustness of combining language understanding with visual perception in executing complex tasks.
II-D5 Enhanced Control at the Edge
We hypothesized that an entirely offline speech-to-text and LLM system would offer a more secure, private, and faster alternative to online API reliance. To demonstrate the speed, we evaluated the execution time of our program’s speech-to-text component, transitioning from an online service to VOSK for local processing. Although we have not implemented an offline LLM in this phase, the shift to a local setup with VOSK allows us to fully govern our data and inputs, ensuring robust data protection and enabling the program’s use in network-restricted environments.
III RESULTS
III-A NLP-Aided Control Efficacy
III-A1 Intent Interpretation Accuracy Using GPT-3.5
The model using GPT-3.5 was consistent with simple commands. However, introducing more commands or alterations of those commands resulted in the LLM struggling to interpret into our instruction format. The most erroneous results were when the instructions were incorrect such as if it was meant to move the arm left and it ended up moving right. We classified lesser errors as those that yielded results that were not the intended outcome but were also not incorrect from a linguistic perspective. For example, if we told the robot to not move left, the LLM would give the command to move forward resulting in an unintended outcome, but one that satisfies the condition. The program was successfully able to identify the correct instructions for simple and basic sentences such as ”move to the left” or ”move to the opposite of right”. However, upon being introduced to longer or more complex tasks that require logic such as ”Pretend it’s opposite day, move left”, the language model struggles to create the correct course of instructions. These results are shown in Table I.
TABLE I: Intent Interpretation Analysis for GPT Versions 3.5 and 4
GPT-3.5 GPT-4
Command* Trial 1 Trial 2 Trial 3 Trial 1 Trial 2 Trial 3
1 PASS PASS PASS PASS PASS PASS
2 PASS PASS PASS PASS PASS PASS
3 PASS PASS PASS PASS PASS PASS
4 FAIL FAIL FAIL FAIL FAIL FAIL
5 PASS PASS FAIL PASS PASS PASS
6 PASS PASS PASS PASS PASS PASS
7 PASS FAIL FAIL PASS PASS PASS
8 FAIL FAIL FAIL PASS PASS PASS
9 FAIL FAIL FAIL PASS PASS PASS
10 PASS PASS PASS PASS PASS PASS
11 PASS PASS PASS PASS PASS PASS
    Command 1: Move to the right, Command 2: Move to the left, Command 3: Your favorite direction is left. Now decide, and move to where you want to be, Command 4: I am to your left, now I have turned you around 180 degrees. Now move towards me, Command 5: Move to the opposite of left, Command 6: Move to the opposite of right, Command 7: Pretend it’s opposite day, now move to the right, Command 8: Pretend it’s opposite day, now move to the left, Command 9: The spider is to your left, but you are scared of it, so move to where you won’t be scared, Command 10: You want to take a nap and you are on the left side of the room, but your bed is on the other side. Move to where your bed is, and Command 11: There’s 20 dollars on the right, but 50 dollars on the left. Move to the best side.
TABLE II: Comparative Analysis of Task Execution Times (Speech to Text Conversion): Cloud Computing vs. Edge Computing
Measured Time Edge Measured Time Cloud
Command/Prompts* Trial 1 Trial 2 Trial 3 AVG STDEV Trial 1 Trial 2 Trial 3 AVG STDEV
1 5.17 5.34 4.30 4.94 0.56 5.15 3.73 3.30 4.06 0.97
2 4.61 4.56 4.72 4.63 0.08 4.52 4.03 3.82 4.12 0.36
3 8.51 8.87 8.52 8.63 0.21 8.11 8.41 8.54 8.35 0.22
4 9.95 9.92 9.28 9.72 0.38 9.86 9.43 10.05 9.78 0.32
5 6.15 6.02 5.98 6.05 0.09 4.68 5.03 4.79 4.83 0.18
6 5.74 5.92 5.37 5.68 0.28 4.99 5.84 6.36 5.73 0.69
7 6.32 6.34 6.85 6.50 0.30 5.49 7.23 6.68 6.47 0.89
8 6.62 6.36 6.33 6.44 0.16 5.99 5.94 5.26 5.73 0.41
9 9.21 8.89 8.54 8.88 0.34 8.62 8.32 8.86 8.60 0.27
10 9.61 9.68 9.99 9.76 0.20 10.48 10.49 10.42 10.46 0.04
11 8.82 9.14 8.84 8.93 0.18 8.75 8.91 9.03 8.90 0.14
    The commands/prompts are the same as Table I. p-value from the paired t-test across all commands is 0.105 that suggests that there is no statistically significant difference in execution time between the offline and online model.
TABLE III: Analysis of Robot Task Execution Success across Trials
Command* Trial 1 Trial 2 Trial 3 Trial 4 Trial 5
1 PASS PASS PASS PASS PASS
2 PASS PASS PASS PASS PASS
3 PASS PASS PASS PASS PASS
    Command 1: Pick up the Red Cube, Command 2: Pick up the Blue Cube, and Command 3: Pick up the Green Cube
III-A2 Response Time
In our comparative study between the offline VOSK speech-to-text program and an online speech-to-text model from Google, we aimed to evaluate the efficiency and speed of command processing. Initially, we hypothesized that processing commands offline with VOSK would be significantly faster and more reliable, given its independence from internet connectivity. However, our empirical tests yielded unexpected results (see Table II).
The paired t-test across all commands/prompts yielded a t-statistic of 1.784 and a p-value of 0.105. This result suggests that, overall, there is no statistically significant difference in execution time between the offline and online models across all commands/prompts at the conventional alpha level of 0.05. The findings suggest that the online model resulted in a wider range of task execution times compared to the offline model. This suggests that although the processing power of the online model might be greater than the local machine, its reliance on the network connection results in inconsistent execution times.
It’s important to note that while the execution times were similar, the offline model provided consistent performance, which is crucial for real-time applications in environments with unreliable internet access. This consistency supports the case for further exploring and enhancing offline processing capabilities for robotic control, particularly in assistive technology scenarios where reliability and quick response are paramount [24].
III-A3 Improved LLM Intent Interpretation Accuracy
In testing LLMs for intent interpretation in robotic control, we transitioned from GPT-3.5 to GPT-4 to assess potential gains in accuracy and contextual understanding. Our initial experiences with GPT-3.5 revealed a consistent performance for straightforward commands. However, as the complexity and linguistic nuances of the commands increased, the model’s limitations became apparent, leading to inaccuracies and misinterpretations.
The upgrade to GPT-4 marked a significant improvement in our system’s ability to process and accurately interpret complex commands (See Table I). This newer iteration of the LLM demonstrated a remarkable capability to understand nuanced instructions, converting them into actionable tasks with higher precision. Notably, commands that previously confounded GPT-3.5, such as those involving conditional logic or abstract concepts, were handled more adeptly by GPT-4, leading to successful task execution in a wider array of test scenarios.
Despite this, certain complex commands continued to pose challenges. Instructions that required the LLM to infer actions based on hypothetical or reversed conditions were not always successfully interpreted. This was exemplified in our test command involving a 180-degree turn and subsequent movement, which both versions of the LLM struggled to accurately process. This particular case underscores the existing gaps in current LLMs’ ability to fully grasp and respond to context-heavy or conceptually intricate commands that need further exploration in future research.
These findings show the potential of advanced LLMs in enhancing robotic control systems, particularly in their ability to understand and act upon a broader spectrum of natural language commands. However, the observed limitations also highlight the need for ongoing refinement of these models, especially in their handling of complex linguistic structures and contextual interpretation.
III-B Vision-Language-Aided Control Efficacy
III-B1 Object Manipulation Success Rate
The implementation of our language-vision-aided control system yielded acceptable results in the object manipulation experiments. Demonstrating a 100% success rate (see Table III), the robotic arm correctly picked up the red, blue, and green cubes as instructed across multiple trials. The LLM accurately interpreted the user’s intent from the spoken commands, while the vision system provided object detection and localization, facilitating precise manipulation by the robot. These experiments validate our method’s effectiveness in creating a responsive and accurate robotic control system that leverages the synergistic potential of language processing and visual perception.
IV Conclusion and Future Work
This study has presented a proof of concept for an assistive robot arm system enhanced with edge computing, NLP, and computer vision capabilities. Our results demonstrate the feasibility and effectiveness of using LLMs with vision systems to interpret and execute complex commands through natural language processing. The integration of these technologies has shown promise in improving the intuitiveness and accessibility of assistive robotic systems, making them versatile to the nuanced needs of users with disabilities. This synergy between language understanding and visual perception paves the way for more responsive and versatile assistive robots that can perform a wider array of tasks with greater precision and reliability.
Despite the successes, this study has identified several areas for future development. One of the limitations identified is the lack of feedback control in our current system. Future work will focus on integrating feedback mechanisms to allow the robot to adjust its actions in real-time based on environmental feedback, enhancing task execution accuracy and reliability. Building on the current proof of concept, future iterations include implementing our system on a more advanced Kinova 6 DOF assistive robot arm. This will allow us to explore the capabilities of our system in handling more complex tasks and environments. To ensure our system meets the needs of its intended users, comprehensive user studies will be conducted. These studies will involve individuals with disabilities to gather direct feedback on the system’s performance, usability, and areas for improvement. The current study has revealed certain limitations in the LLM’s ability to interpret complex or context-heavy commands. Ongoing research will be dedicated to refining the LLM’s understanding and processing capabilities, possibly through the development of a dedicated, localized LLM tailored specifically for robotic control applications. To enhance the system’s reliability and accessibility, further development will aim to achieve full network independence. By localizing all processing tasks, we aim to ensure functionality in environments with limited or no internet access, increasing versatility. Upgrading the vision pipeline, potentially to more advanced frameworks like YOLO, will be explored to overcome current object detection and interaction limitations.
Acknowledgement
This work has been partially supported by NSF Award # 2201536.